A survey classified 200 students by gender and by their opinion on a certain issue. The number falling into the different categories are shown in the following table. A student is randomly chosen from the group.
Opinion
Gender For Against Total
Male 30 40 70
Female 50 80 130
Total 80 120 200

9. The probability that the student is female and is against the issue is

10. The probability that the student is male or is for the issue is

11. Given that the student chosen is for the issue, the conditional probability that the
student is male is
1
Expert's answer
2020-07-21T11:12:26-0400

9. The probability that the student is female

and is against the issue =80/200=0.4

10. The probability that the student is male

or is for the issue =(70+80-30)/200=120/200=0.6

11. Given that the student chosen is for the issue,

the conditional probability that the student is male

=(30/200)/(80/200)=3/8